OpenStudy (anonymous):

How to find a one-family parameter of solutions of the differential equation xy' = y? Verify that each member of the family is a solution of the initial-value problem xy' = y, y(0) = 0. How do I approach these kind of problems?

OpenStudy (anonymous):

By inspection* find a one-family parameter of solutions...

OpenStudy (abb0t):

First, get it in it's general form.

OpenStudy (anonymous):

Manipulating the expression we find \(y'/y=1/x\) and we know \(y'/y=(\log y)'\) ergo:$$(\log y)'=1/x\\\log y=\int\frac1x\ dx=\log x+C\\y=kx$$where \(k=e^C\)

OpenStudy (anonymous):

It is a separable equation.

OpenStudy (anonymous):

Alternatively you can just recall that \(y'\) is effectively the slope of \(y\); when we write \(y=y'x\) it suggests that this function has constant slope :-)

OpenStudy (anonymous):

hmm:$$y'=y/x\\\frac{dy}{dx}=\frac{y}x$$does this make it more obvious that the function must be a line thru the origin? :-)

OpenStudy (anonymous):

Yes it does. Thank you. :)